\begin{thebibliography}{ZZZZZZ}

\bibitem[BBGN]{BBGN} V. Balaji, I. Biswas, O. Gabber and D.
S. Nagaraj, Brauer obstruction for a universal vector bundle,
{\it Comp. Ren. Acad. Sci. Math.} \textbf{345} (2007), 265--268.

\bibitem[BH]{BH} I. Biswas and Y. I. Holla, Brauer group of moduli of principal 
bundles over a curve, {\it Jour. Reine Ang. Math.} \textbf{677} (2013), 225--249.

\bibitem[dJ]{dJ} A.J. de Jong, A result of Gabber,\\
http://www.math.columbia.edu/$\sim$dejong/papers/2-gabber.pdf.

\bibitem[DN]{DN} J.-M. Drezet and M. S. Narasimhan, Groupe de Picard
des vari\'et\'es de modules de fibr\'es semi-stables sur les
courbes alg\'ebriques, {\it Invent. Math.} {\bf 97} (1989), 53--94.

\bibitem[Ga]{Ga} O. Gabber, Some theorems on Azumaya algebras, in {\it The Brauer group}
(Sem., Les Plans-sur-Bex, 1980), pp. 129--209, Lecture Notes in Math., 844, Springer,
Berlin-New York, 1981.

\bibitem[Gr1]{Gr} A. Grothendieck, Le groupe de Brauer, I,
in \textit{Dix expos\'es sur la cohomologie des sch\'emas}, pp. 67--87,
Advanced Studies in Pure Mathematics, 3, North-Holland, Amsterdam, 1968.

\bibitem[Gr2]{Gr2} A. Grothendieck, Le groupe de Brauer, I,
in \textit{Dix expos\'es sur la cohomologie des sch\'emas}, pp. 67--87,
Advanced Studies in Pure Mathematics, 3, North-Holland, Amsterdam, 1968.

\bibitem[Gr3]{Gr3} A. Grothendieck, Le groupe de Brauer, I,
in \textit{Dix expos\'es sur la cohomologie des sch\'emas}, pp. 88--188,
Advanced Studies in Pure Mathematics, 3, North-Holland, Amsterdam, 1968.

\bibitem[Hof]{Ho} N. Hoffmann, The Picard group of a coarse moduli space of
vector bundles in positive characteristic, {\it Cent. Eur. Jour. Math.}
{\bf 10} (2012), 1306--1313.

\bibitem[Hoo]{Hoo} R. T. Hoobler, When is ${\rm Br}(X)\,=\,{\rm Br}'(X)$?
in {\it Brauer groups in ring theory and algebraic geometry} (Wilrijk, 1981), pp.
231--244, Lecture Notes in Math., 917, Springer, Berlin-New York, 1982.

\bibitem[Mi]{Mi} J. S. Milne, \textit{\'Etale cohomology}, Princeton Mathematical
Series, 33, Princeton University Press, 1980.

\bibitem[Mu]{Mu} D. Mumford, \textit{Abelian varieties}. With appendices by C. P. 
Ramanujam and Yuri Manin. Corrected reprint of the second (1974) edition. Tata 
Institute of Fundamental Research Studies in Mathematics, 5. Published for the Tata 
Institute of Fundamental Research, Bombay; by Hindustan Book Agency, New Delhi, 
2008.

\bibitem[Ne]{Ne} P. E. Newstead, {\it Introduction to moduli problems and orbit spaces},
Tata Institute of Fundamental Research Lectures on Mathematics and Physics, 51,
Narosa Publishing House, New Delhi, 1978.

\bibitem[Ra]{Ra} S. Ramanan, The moduli spaces of vector bundles over an algebraic 
curve, \textit{Math. Ann.} \textbf{200} (1973), 69--84.

\end{thebibliography}

\documentclass[12pt,reqno]{amsart}

\usepackage{amssymb}

\newtheorem{theorem}{Theorem}[section]
\newtheorem{proposition}[theorem]{Proposition}
\newtheorem{lemma}[theorem]{Lemma}

\numberwithin{equation}{section}

\begin{document}

\baselineskip=15.5pt

\title[Brauer group of the moduli spaces of stable vector bundles]{Brauer group of
the moduli spaces of stable vector bundles of fixed determinant over a smooth curve}

\author[I. Biswas]{Indranil Biswas}

\address{School of Mathematics, Tata Institute of Fundamental
Research, Homi Bhabha Road, Mumbai 400005, India}

\email{indranil@math.tifr.res.in}

\author[T. Sengupta]{Tathagata Sengupta}

\address{Department of Mathematics, University of Hyderabad, C. R. Rao Road,
Hyderabad 500046, India}

\email{tsengupta@gmail.com}

\subjclass[2010]{14H60, 14F22, 14D23, 53C08}

\keywords{Moduli space, Brauer group, $\mu_r$-gerbe, Picard group.}

\date{}

\begin{abstract}
Let $X$ be an irreducible smooth projective curve, defined over an algebraically closed 
field $k$, of genus at least three and $L$ a line bundle on $X$. Let ${\mathcal 
M}_X(r,L)$ be the moduli space of stable vector bundles on $X$ of rank $r$ and 
determinant $L$ with $r\, \geq\, 2$. We prove that the Brauer group ${\rm 
Br}(\mathcal{M}_X(r,L))$ is cyclic of order ${\rm g.c.d.}(r,{\rm degree}(L))$. We also 
prove that ${\rm Br}(\mathcal{M}_X(r,L))$ is generated by the class of the projective 
bundle obtained by restricting the universal projective bundle. These results were 
proved earlier in \cite{BBGN} under the assumption that $k\,=\,\mathbb C$.
\end{abstract}

\maketitle

\section{Introduction}

Let $X$ be a compact connected Riemann surface of genus $g$, with $g\, \geq\, 3$. 
Fix a holomorphic line bundle $L$ over $X$ and also fix an integer $r\, \geq\, 2$. 
Let ${\mathcal M}_X(r,L)$ denote the moduli space of stable vector bundles on $X$ of 
rank $r$ and determinant $L$, which is a smooth quasiprojective complex variety of 
dimension $(r^2-1)(g-1)$. There is a Poincar\'e vector bundle over $X\times 
{\mathcal M}_X(r,L)$ if and only if $r$ and $\text{degree}(L)$ are coprime 
\cite{Ra}. When $r$ and $\text{degree}(L)$ are coprime, any two Poincar\'e vector 
bundle over $X\times {\mathcal M}_X(r,L)$ differ by tensoring with a line bundle 
pulled back from ${\mathcal M}_X(r,L)$. Hence the projectivized Poincar\'e bundle in 
unique. Even when $r$ and $\text{degree}(L)$ are not coprime, there is a unique 
projective Poincar\'e bundle over $X\times {\mathcal M}_X(r,L)$, although it is not 
a projectivization of a vector bundle.

In \cite{BBGN} it was proved that the Brauer group of ${\mathcal M}_X(r,L)$ is 
cyclic of order ${\rm g.c.d.}(r,{\rm degree}(L))$. As mentioned above, there is a 
universal projective bundle $\mathcal P$ on $X\times {\mathcal M}_X(r,L)$. Fixing a 
point $x\, \in\, X$, let ${\mathcal P}_x$ be the projective bundle on ${\mathcal 
M}_X(r,L)$ obtained by restricting $\mathcal P$ to $\{x\}\times {\mathcal 
M}_X(r,L)$. In \cite{BBGN} it was also shown that the Brauer group 
$\text{Br}({\mathcal M}_X(r,L))$ is generated by the class of ${\mathcal P}_x$.

Our aim here is to prove these results for all algebraically closed fields;
see Theorem \ref{thm1}.

The computation in \cite{BBGN} crucially uses the calculation of the Picard group of 
${\mathcal M}_X(r,L)$. It may be mentioned that the assumption in \cite{DN} that the 
characteristic of the base field is zero is used in the computation of the Picard 
group of the moduli space ${\mathcal M}_X(r,L)$. In particular, the Reynolds' 
operators, which play a crucial role in the computation, are valid only in 
characteristic zero. A recent theorem of Hoffmann shows that the Picard group of the 
moduli space does not depend on the base field \cite{Ho}. The proof of
Theorem \ref{thm1} follows the strategy of \cite{BBGN}; some details not
given in \cite{BBGN} are given here.

\section{Universal projective bundle and Brauer group}

Let $k$ be an algebraically closed field. Let $X$ be an irreducible smooth
projective curve, defined over $k$, of genus $g$, with $g\, \geq\, 3$. Fix
an integer $r\, \geq\, 2$ and also fix a line bundle $L$ over $X$. The degree
of $L$ will be denoted by $d$. Let ${\mathcal M}_X(r,d)$ be the
moduli space of stable vector bundles on $X$ of rank $r$ and degree $d$.
Consider the morphism
$$
\phi\, :\, {\mathcal M}_X(r,d)\, \longrightarrow\, \text{Pic}^d(X)\, ,
\ \ \ E\, \longmapsto\, \bigwedge\nolimits^r E\, .
$$
Let
$$
{\mathcal M}_X\,=\, {\mathcal M}_X(r,L)\, :=\, \phi^{-1}(L)
$$
be the fiber of $\phi$ over the point $L\, \in\, \text{Pic}^d(X)$. This moduli
space ${\mathcal M}_X$ is canonically identified with the following two moduli
spaces:
\begin{enumerate}
\item moduli space of pairs of the form $(E,\, \xi)$, where $E$ is a
stable vector bundle over $X$ of rank $r$, and $\xi\, :\, \bigwedge^r E
\, \longrightarrow\, L$ is an isomorphism, and

\item the moduli space of stable vector bundles $E$ on $X$ of rank $r$ such that
$\bigwedge^r E$ is isomorphic to $L$
\end{enumerate}
(see \cite[p.~1308, Proposition 2.1]{Ho}).

It is known that there is a universal projective bundle
\begin{equation}\label{bp}
{\mathcal P}\, \longrightarrow\, X\times {\mathcal M}_X
\end{equation}
(\cite{Ra}, \cite{Ne}).
This follows from the construction of the moduli space and the fact that
the global automorphisms of a stable vector bundle are nonzero constant
scalar multiplications; this projective bundle $\mathcal P$ is described in the proof of
Proposition \ref{prop1}. Fix a closed point $x\, \in\, X$. Let
\begin{equation}\label{e1}
{\mathcal P}_x\,:=\, {\mathcal P}\vert_{\{x\}\times {\mathcal M}_X}
\, \stackrel{f}{\longrightarrow}\, {\mathcal M}_X
\end{equation}
be the restriction of ${\mathcal P}$ to $\{x\}\times {\mathcal M}_X$.

For any quasiprojective variety $Y$ defined over the field $k$, the Brauer group 
$\text{Br}(Y)$ of $Y$ is defined to be the Morita equivalence classes of Azumaya 
algebras over the variety $Y$. It is known that this Brauer group $\text{Br}(Y)$ 
coincides with the equivalence classes of all principal $\text{PGL}_k$--bundles over 
$Y$, where two principal $\text{PGL}_k$--bundles $P$ and $Q$ are identified if there 
are two vector bundles $V_1$ and $V_2$ over $Y$ satisfying the condition that the two 
principal $\text{PGL}_k$--bundles $P\otimes {\mathbb P}(V_1)$ and $Q\otimes {\mathbb 
P}(V_2)$ are isomorphic. The addition of two projective bundles $P$ and $Q$ in the
Brauer group $\text{Br}(Y)$ is defined to be the equivalence class of the projective bundle
$P\otimes Q$. The inverse of a projective bundle $P$ in $\text{Br}(Y)$ is the
equivalence class of the dual projective bundle $P^*$. (See \cite{Gr}, \cite{Gr2},
\cite{Gr3}, \cite{Mi}, \cite{Ga} for properties of Brauer groups.)
The cohomological Brauer group $\text{Br}'(Y)$ of the variety 
$Y$ is the torsion part of the \'etale cohomology group $H^2_{et}(Y,\, {\mathbb G}_m)$. 
There is a natural injective homomorphism $\text{Br}(Y)\,\longrightarrow\, 
\text{Br}'(Y)$ which is in fact an isomorphism by a theorem of Gabber \cite{dJ}, 
\cite{Hoo}.

\begin{proposition}\label{prop1}
The Brauer group ${\rm Br}(\mathcal{M}_X)$ is generated by the
class $cl({\mathcal P}_x)\, \in\, {\rm Br}(\mathcal{M}_X)$ of the projective
bundle ${\mathcal P}_x$ defined in \eqref{e1}.
\end{proposition}

\begin{proof}
Given any line bundle $L_0$ on $X$, the morphism
$$
{\mathcal M}_X\,=\, {\mathcal M}_X(r,L)\,\longrightarrow\,
{\mathcal M}_X(r,L\otimes L^r_0)\, ,\ \ E\, \longmapsto\, E\otimes L_0
$$
is an isomorphism. The natural isomorphism of ${\mathbb P}(E\otimes L_0)$ with
${\mathbb P}(E)$ produces an isomorphism between the universal projective bundles
over $X\times {\mathcal M}_X(r,L)$ and $X\times {\mathcal M}_X(r,L\otimes L^r_0)$.
Therefore, after tensoring with a line bundle $L_0$ of sufficiently large
degree, we may assume that
$$
\frac{d}{r} \,>\, 2g-1\, .
$$
Let $\overline{\mathcal M}_X$ denote the moduli space of semistable vector bundles
$E$ on $X$ of rank $r$ with $\bigwedge^r E\,=\, L$.

The cotangent bundle of $X$ will be denoted by $K_X$.
For any vector bundle $E\, \in\, \overline{\mathcal M}_X$ and any point $y\, \in\, X$,
$$
H^1(Y,\, E\otimes {\mathcal O}_X(-y))\,=\,H^0(Y,\, E^*\otimes K_X
\otimes {\mathcal O}_X(y))^*\,=\, 0
$$
because ${\rm degree}(E^*\otimes K_X\otimes{\mathcal O}_X(y))\, <\, 0$
and $E^*\otimes K_X\otimes {\mathcal O}_X(y)$ is semistable. So from the
long exact sequence of cohomologies associated to the short exact sequence
$$
0\,\longrightarrow\,E\otimes {\mathcal O}_X(-y)\,\longrightarrow\, E
\,\longrightarrow\, E_y \,\longrightarrow\, 0
$$
it follows that the evaluation homomorphism $H^0(X,\, E)\,\longrightarrow\, E_y$
is surjective; hence $E$ is generated by its global sections.

Take any $E\,\in\, \overline{\mathcal M}_X$. Since the vector bundle $E$ is
generated by its global sections, there is a short exact sequence 
\begin{equation}\label{req}
0\,\longrightarrow\, \mathcal{O}^{\oplus (r-1)}_X\,\longrightarrow\, E\,
\longrightarrow\,\bigwedge\nolimits^r E\,=\,L\,\longrightarrow\, 0\, .
\end{equation}
This short exact sequence
does not split because $E$ is semistable and $\text{degree}(L)\, >\,0$.
All such nontrivial extensions are parameterized by
$$\mathbb{P}(H^1(X, {\rm Hom}(L,\mathcal{O}^{\oplus(r-1)}_X))^*)\,=\,
\mathbb{P}((H^1(X,L^*)^*\otimes_k k^{\oplus (r-1)}) \,=\, \mathbb{P}((H^1(X,L^*)^{
\oplus (r-1)})^*)\, .$$
The standard action of $\text{GL}(r-1,{k})$ on $k^{\oplus (r-1)}$ produces an action of
$\text{GL}(r-1,{k})$ on the projective space $\mathbb{P}((H^1(X,L^*)^{\oplus (r-1)})^*)$. The
moduli space $\overline{\mathcal M}_X$ is the geometric invariant theoretic quotient
$$
\mathbb{P}((H^1(X,L^*)^{r-1})^*)/\!\!/\text{GL}(r-1,{k})\,=\,
\mathbb{P}((H^1(X,L^*)^{r-1})^*)/\!\!/\text{PGL}(r-1,{k})\,=\,
\overline{\mathcal M}_X
$$
(see \cite{Ne}, \cite{DN}).

The tautological line bundle ${\mathcal O}_{\mathbb{P}((H^1(X,L^*)^{r-1})^*)}(1)$
on $\mathbb{P}((H^1(X,L^*)^{r-1})^*)$ will be denoted by ${\mathcal L}_0$. Let
$$
p_1\, :\, X\times\mathbb{P}((H^1(X,L^*)^{r-1})^*)\,\longrightarrow\,X\, ,\ \
p_2\, :\, X\times\mathbb{P}((H^1(X,L^*)^{r-1})^*)\,\longrightarrow\,
\mathbb{P}((H^1(X,L^*)^{r-1})^*)
$$
be the natural projections. There is a universal extension over
$X\times\mathbb{P}((H^1(X,L^*)^{r-1})^*)$
\begin{equation}\label{a1}
0 \,\longrightarrow\, (p^*_1 \mathcal{O}^{r-1}_X)\otimes p^*_2{\mathcal L}_0
\,\longrightarrow\, {\mathcal E} \,\longrightarrow\, p^*_1 L\,\longrightarrow\, 0\, .
\end{equation}
Let ${\mathcal U}\, \subset\, \mathbb{P}((H^1(X,L^*)^{r-1})^*)$ be the subset
defined by all points
$t\, \in\, \mathbb{P}((H^1(X,L^*)^{r-1})^*)$ such that the vector bundle ${\mathcal
E}\vert_{X\times\{t\}}$ is stable. This subset $\mathcal U$ is nonempty Zariski open. Let
\begin{equation}\label{t}
\theta\,:\, {\mathcal U}\,\longrightarrow\,{\mathcal U}/\!\!/ \text{PGL}(r-1,{k})\,=\,
{\mathcal M}_X
\end{equation}
be the quotient map. Consider the action of
$\text{PGL}(r-1,{k})$ on $X\times \mathbb{P}((H^1(X,L^*)^{r-1})^*)$ given by the
trivial action on $X$ and the above action of $\mathbb{P}((H^1(X,L^*)^{r-1})^*)$.
This action lifts to an action of
$\text{PGL}(r-1,{k})$ on ${\mathbb P}({\mathcal E})$. The corresponding
geometric invariant theoretic quotient
$$
{\mathcal P}\, :=\, 
({\mathbb P}({\mathcal E})\vert_{X\times{\mathcal U}})/\!\!/\text{PGL}(r-1,{k})
$$
is the universal projective bundle on $X\times {\mathcal M}_X$ (see \eqref{bp}).

Consider the map
\begin{equation}\label{F}
F\,:=\, \text{Id}_X\times f\, :\, X\times {\mathcal P}_x\,\longrightarrow\,
X\times{\mathcal M}_X\, ,
\end{equation}
where $f$ is the projection in \eqref{e1}. We will construct a vector bundle
$$
{\mathcal V} \,\longrightarrow\, X\times {\mathcal P}_x
$$
with the property that ${\mathbb P}({\mathcal V})\,=\, F^*{\mathcal P}$.

Let ${\mathcal E}_x\,:=\, {\mathcal E}\vert_{\{x\}\times {\mathcal 
U}}\,\longrightarrow\, {\mathcal U}$ be the vector bundle obtained by restricting 
$\mathcal E$ in \eqref{a1} to $\{x\}\times {\mathcal U}$. Let
$$
{\mathcal Q}\,:=\, {\mathbb 
P}({\mathcal E}_x)\, \stackrel{\beta'}{\longrightarrow}\, {\mathcal U}
$$ be the 
corresponding projective bundle. Define
$$
\beta\, :=\, \text{Id}_X\times\beta'\, :\, 
X\times {\mathcal Q} \,\longrightarrow\, X\times {\mathcal U}\, ,
$$
and consider the pulled back vector bundle
$$
\widetilde{\mathcal{E}}\,:=\, (\beta^*\mathcal{E})\otimes (q^*_2
{\mathcal O}_{\mathcal Q}(-1)) \,\longrightarrow\,X\times {\mathcal Q}\, ,
$$
where $q_2\, :\, X\times {\mathcal Q}\,\longrightarrow\,{\mathcal Q}$ is the natural
projection, and $${\mathcal O}_{\mathcal Q}(1)\,\longrightarrow\,
{\mathbb P}({\mathcal E}_x)\,=\,
{\mathcal Q}$$ is the tautological line bundle. For the natural action of $\text{GL}(r-1,
k)$ on $\widetilde{\mathcal{E}}$, the center ${\mathbb G}_m$ of $\text{GL}(r-1,{k})$
acts trivially on $\widetilde{\mathcal{E}}$. Consequently, the geometric invariant
theoretic quotient
$$
{\mathcal V} \,:=\, \widetilde{\mathcal{E}}/\!\!/\text{GL}(r-1,{k}) \,
{\longrightarrow}
\, X\times({\mathcal Q}/\!\!/\text{GL}(r-1,{k}))\,=\, X\times {\mathcal P}_x
$$
is a vector bundle. It is straight-forward to check that
\begin{itemize}
\item ${\mathbb P}({\mathcal V})\,=\, F^*{\mathcal P}$, where $F$ is the map in
\eqref{F}, and

\item for each point $y\, \in\,{\mathcal P}_x$, the vector bundle
${\mathcal V}\vert_{X\times\{y\}}$ on $X$ lies in the isomorphism class of
vector bundles associated to the point $f(y)\, \in\,{\mathcal M}_X$, where
$f$ is defined in \eqref{e1}.
\end{itemize}

Let $B\, :\, X\times {\mathcal P}_x\,\longrightarrow\,{\mathcal P}_x$
be the natural projection. 
Consider the direct image $B_*{\mathcal V}\, \longrightarrow\, {\mathcal P}_x$.
Let $Z\,\subset\, {\mathcal P}_x$ be a nonempty Zariski open subset
such that the restriction $(B_*{\mathcal V})\vert_Z$ is a trivial vector bundle.
Fix a trivialization of $(B_*{\mathcal V})\vert_Z$. Take a point $y_0\, \in\,
Z$ and choose $r-1$ linearly independent sections $$s_1,\, \cdots ,\, s_{r-1}\, \in\,
H^0(X\times\{y_0\},\, {\mathcal V}\vert_{X\times\{y_0\}})$$
such that the coherent subsheaf of ${\mathcal V}\vert_{X\times\{y_0\}}$ generated by
$s_1,\, \cdots ,\, s_{r-1}$ is a subbundle of ${\mathcal V}\vert_{X\times\{y_0\}}$
of rank $r-1$; we note that from \eqref{req} it follows immediately that such
$r-1$ linearly independent sections exist. Extend each $s_i$ to a section $\widetilde{s}_i$
of ${\mathcal V}\vert_{X\times Z}$ using the above trivialization of
$(B_*{\mathcal V})\vert_Z$. There is a Zariski open subset $Z'\, \subset\, Z$ containing
$y_0$ such that the coherent subsheaf of ${\mathcal V}\vert_{X\times Z}$ generated by
$\widetilde{s}_1,\, \cdots ,\, \widetilde{s}_{r-1}$ is a subbundle of
${\mathcal V}\vert_{X\times Z'}$ of rank $r-1$. Note that this subbundle
over $X\times Z'$ is trivial and a trivialization is given by the images
of $\widetilde{s}_1,\, \cdots ,\, \widetilde{s}_{r-1}$. Therefore, on $X\times Z'$, we have a
short exact sequence of vector bundles
$$
0 \,\longrightarrow\, {\mathcal O}^{\oplus (r-1)}_{X\times Z'}
\,\longrightarrow\, {\mathcal V}\vert_{X\times Z'} \,\longrightarrow\, L'
\,\longrightarrow\, 0\, ,
$$
where $L'$ is a line bundle on $X\times Z'$. Considering the top exterior products
it follows that for each point $y\, \in\, Z'$,
the restriction $L'\vert_{X\times\{y\}}$ is isomorphic to the line bundle $L$.
Now from the seesaw theorem (see \cite[p.~51, Corollary~6]{Mu}) it follows that
there is a line bundle $L''$ on $Z'$ such that the line bundle $L'\otimes B^*L''$ 
on $X\times Z'$ is isomorphic to the pullback of $L$ to $X\times Z'$. We may trivialize
$L''$ over suitable nonempty Zariski open subsets of $Z'$. Therefore, it follows 
that there is a nonempty Zariski open subset
\begin{equation}\label{i}
\iota\, :\, {\mathcal W}\, \hookrightarrow\, Z' \, \subset\, {\mathcal P}_x
\end{equation}
such that the restriction $L'\vert_{X\times\mathcal W}$ is isomorphic to
the pullback of $L$ to $X\times\mathcal W$.

Consequently, there is a morphism
$$
\varphi\, :\, {\mathcal W}\,\longrightarrow\, {\mathcal U}\, \subset\,
\mathbb{P}((H^1(X,L^*)^{r-1})^*)
$$
such that the following diagram is commutative
\begin{equation}\label{c}
\begin{matrix}
{\mathcal W}& \stackrel{\varphi}{\longrightarrow}& {\mathcal U}\\
~ \Big\downarrow \iota && ~\Big\downarrow\theta \\
{\mathcal P}_x& \stackrel{f}{\longrightarrow}& {\mathcal M}_X
\end{matrix}
\end{equation}
where $\theta$, $\iota$ and $f$ are the morphisms in \eqref{t}, \eqref{i} and
\eqref{e1} respectively.

The codimension of the complement $$\mathbb{P}((H^1(X,L^*)^{r-1})^*)\setminus
{\mathcal U}\, \subset\,\mathbb{P}((H^1(X,L^*)^{r-1})^*)$$ is at least two.
To prove this, note that $\text{Pic}({\mathcal U})\,=\, \mathbb Z$ \cite[p.~89,
Proposition 7.13]{DN} (here we need the assumption that $g\, \geq\, 3$); this immediately
implies that the codimension of the complement ${\mathcal U}^c$ is at least two. Since the
Brauer group of a projective space is zero, in view of this codimension estimate,
it follows from the ``Cohomological purity'' \cite[p.~241, Theorem~VI.5.1]{Mi} (it also
follows from \cite[p.~292--293]{Gr}) that
\begin{equation}\label{bz}
{\rm Br}({\mathcal U})\, =\, 0\, .
\end{equation}
{}From the commutativity of \eqref{c} we conclude that the pullback homomorphism
$$
\iota^*\circ f^* \, =\, (f\circ \iota)^* \, :\, \text{Br}({\mathcal M}_X)
\,\longrightarrow\, \text{Br}({\mathcal W})
$$
coincides with the homomorphism $\varphi^*\circ \theta^*\, :\, \text{Br}({\mathcal M}_X) 
\,\longrightarrow\, \text{Br}({\mathcal W})$. On the other hand, from \eqref{bz}
we know that $\varphi^*\circ \theta^*\,=\, 0$. Hence
$$
\iota^*\circ f^* \, =\,0\, .
$$

On the other hand, the homomorphism
$$
\iota^*\, :\, \text{Br}({\mathcal P}_x)
\,\longrightarrow\, \text{Br}({\mathcal W})
$$
is injective because $\mathcal W$ is a Zariski open dense subset of
${\mathcal P}_x$ (see \cite[p.~142, Theorem~2.5]{Mi}). Consequently,
$$
f^* \,: \,\text{Br}({\mathcal M}_X)
\,\longrightarrow\,\text{Br}({\mathcal P}_x)
$$
is the zero homomorphism. On the other hand, the kernel of the above
homomorphism $f^*$ is generated by the class of ${\mathcal P}_x$
\cite[p.~193, Theorem~2]{Ga}. Therefore, we conclude that $\text{Br}({\mathcal M}_X)$ is
generated by the class of ${\mathcal P}_x$. This completes the proof.
\end{proof}

We will denote the integer ${\rm g.c.d.}(r,d)$ by $\delta$.

\begin{lemma}\label{lem1}
The order of the Brauer class $cl({\mathcal P}_x)\, \in\, {\rm Br}(\mathcal{M}_X)$
is $\delta$.
\end{lemma}

\begin{proof}
Let $\widetilde{\mathcal M}_X\,=\, \widetilde{\mathcal M}_X(r,L)$ be the moduli stack
of pairs of the form $(E,\, \xi)$, where $E$ is a
stable vector bundle over $X$ of rank $r$ and $\xi\, :\, \bigwedge^r E
\, \longrightarrow\, L$ is an isomorphism. Let $\mu_r$ denote the kernel of the
homomorphism
$$
{\mathbb G}_m\, \longrightarrow\, {\mathbb G}_m\, ,\ \ z\, \longmapsto\, z^r\, .
$$
The natural morphism
$$
\gamma\, :\, \widetilde{\mathcal M}_X\,\longrightarrow\, {\mathcal M}_X
$$
makes $\widetilde{\mathcal M}_X$ a $\mu_r$--gerbe over ${\mathcal M}_X$. Let
$$
c_0\,\in\, H^2({\mathcal M}_X,\,\mu_r)
$$
be the class of this $\mu_r$--gerbe. The Brauer class
$$
cl({\mathcal P}_x)\, \in\, {\rm Br}(\mathcal{M}_X)\,=\, H^2({\mathcal M}_X,\,
{\mathbb G}_m)
$$
coincides with the image of $c_0$ under the homomorphism
$$
\eta\, :\, H^2({\mathcal M}_X,\,\mu_r)\, \longrightarrow\, H^2({\mathcal M}_X,\,
{\mathbb G}_m)
$$
given by the inclusion of $\mu_r$ in ${\mathbb G}_m$.

There is a short exact sequence
\begin{equation}\label{es1}
0\, \longrightarrow\, \text{Pic}(\mathcal{M}_X)
\, \stackrel{\gamma^*}{\longrightarrow}\, \text{Pic}(\widetilde{\mathcal M}_X)
\, \stackrel{\nu}{\longrightarrow}\, {\rm Hom}(\mu_r,\, {\mathbb G}_m)\,=\,
{\mathbb Z}/r{\mathbb Z} \, \stackrel{\alpha}{\longrightarrow}\,{\rm Br}(\mathcal{M}_X)
\end{equation}
\cite[p.~232, Lemma 4.4]{BH}, where $\nu$ sends a line bundle on the $\mu_r$--gerbe
$\widetilde{\mathcal M}_X$ to the weight associated
to the action of $\mu_r$ on it. The image
$$
\alpha(1)\, \in\, {\rm Br}(\mathcal{M}_X)\,=\, H^2({\mathcal M}_X,\,
{\mathbb G}_m)
$$
coincides with the image of the class of the $\mu_r$--gerbe $\widetilde{\mathcal M}_X$
under the above homomorphism $\eta$. From this it follows that
\begin{equation}\label{a}
\alpha(1)\,=\, cl({\mathcal P}_x)
\end{equation}
because $cl({\mathcal P}_x)$ also coincides with the image of the class of the
$\mu_r$--gerbe $\widetilde{\mathcal M}_X$ under the above homomorphism $\eta$.

{}From \cite[p.~1311, Lemma 3.6]{Ho}, \cite[p.~1311, Lemma 3.3]{Ho}
and \cite[p.~1310, Theorem 3.1]{Ho} it follows that
$$
{\mathbb Z}/\text{image}(\nu)\,=\, {\mathbb Z}/\delta{\mathbb Z}\, ,
$$
where $\nu$ is the homomorphism in \eqref{es1}. Therefore, from \eqref{es1} we conclude
that the order of $\alpha(1)\, \in\, {\rm Br}(\mathcal{M}_X)$ is $\delta$. Now
the lemma follows from \eqref{a}.
\end{proof}

Combining Proposition \ref{prop1} and Lemma \ref{lem1} we have:

\begin{theorem}\label{thm1}
The Brauer group ${\rm Br}(\mathcal{M}_X)$ is cyclic of order $\delta\,=\,
{\rm g.c.d.}(r,d)$. The group ${\rm Br}(\mathcal{M}_X)$ is generated by the
class $cl({\mathcal P}_x)\, \in\, {\rm Br}(\mathcal{M}_X)$
of the projective bundle ${\mathcal P}_x$ defined in \eqref{e1}.
\end{theorem}

\section*{Acknowledgements}

We thank the referee for helpful comments. The first-named author is supported by a 
J. C. Bose Fellowship. The second-named author would like to thank Tata Institute of 
Fundamental Research for its hospitality while the paper was being worked on.

%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%

\end{document}